OpenStudy (anonymous):

x equals the square root of 8 x

OpenStudy (watchmath):

\(x=\sqrt{8x}\)\ Square both sides \(x^2=8x\) \(x^2-8x=0\) \(x(x-8)=0\) \(x=0,8\)

OpenStudy (anonymous):

X^x = 8x X = 8

OpenStudy (anonymous):

Sorry that's x^2 = 8x X = 8

OpenStudy (anonymous):

um watch math can you do it if it has an equation on both sides or is it different

OpenStudy (anonymous):

See if x^2 = 8x then X*x = 8*x Divide both sides by x X=8

OpenStudy (watchmath):

In this problem we need to bring everything to one side of the equation so that we can factor it. Otherwise if you divide by \(x\) like ishwan did you will loose one solution, namely \(x=0\).

OpenStudy (anonymous):

an example is n-4= the spuare root of 12-3n

OpenStudy (watchmath):

The idea is still the same. You square both sides and collect everything on one side. \((n-4)^2=12-3n\) \(n^2-8n+16=12-3n\) \(n^2-5n+4=0\) \((n-4)(n-1)=0\) \(n=4\) or \(n=1\) Double check to the original equation shows that \(n=1\) is not a solution. Thus \(n=4\) is the only solution. (press good answer button if my answer is helpful :D )

OpenStudy (anonymous):

Watch math is correct

OpenStudy (anonymous):

alright thanks i need help with one more the square root of b-1-b=negative seven

OpenStudy (watchmath):

Post as a new question :)

OpenStudy (anonymous):

the square root of b-1-b equals negative seven

OpenStudy (anonymous):

shawna can paint a fence in 5 hrs. Jassie can paint the same fence in 10 hrs. how long will it take them to paint the fence together